# Joint Communication and Trajectory Design for Movable Antenna Systems

### 05 May 2026 v1

@@ III-A Two-Path Case
$\left|h_{u}[k]\right|^{2}=\left|a_{1}\right|^{2}+\left|a_{2}\right|^{2}+2\left|a_{1}\right|\left|a_{2}\right|\cos\varphi[k],$ (11)
@@ III-B Multi-Path Case
${W}_{i,j}=-\log_{2}\left(1+\frac{P_{t}\left|h_{u}\left(p_{i}\right)\right|^{2}}{\sigma^{2}}\right),\forall(i,j)\in\tilde{E}$. (15)
Based on the above, it can be shown that $\left(P1\right)$ is equivalent to the following fixed-hop shortest path problem $\left(P2\right)$.
@@ V Conclusion
Based on (16), the average rate $\frac{1}{K}\sum_{k=1}^{K}R_{u}[k]$ can be decomposed as $\frac{1}{K}\left(\sum_{k=1}^{k_{1}}R_{u}[k]+(K-k_{1})R_{u}[K]\right)$. Consider an arbitrary feasible trajectory, denoted as $x^{\prime}[k],k\in{\cal K}$, and let $R^{\prime}_{u}[k]$ denote the achievable rate at the receiver for $x=x^{\prime}[k]$. Evidently, we have $(K-k_{1})R_{u}[K]\geq\sum_{k=k_{1}}^{K}R^{\prime}_{u}[k]$. Next, we show $\sum_{k=1}^{k_{1}}R_{u}[k]\geq\sum_{k=1}^{k_{1}}R^{\prime}_{u}[k]$. Note that moving the MA from $x_{0}$ to $x^{\prime}$ within the first to the $k_{1}$-th time slot always yields an increasing channel power gain (see the green curve in Fig. 8), while moving the MA in the opposite direction always yields a decreasing channel power gain first within this period (see the red curve in Fig. 8). Hence, if the trajectory $x^{\prime}[k],k\in{\cal K}$ moves the MA away from $x^{*}$, we have $R_{u}[k]>R^{\prime}_{u}[k],1\leq k\leq k_{1}$, and thus, $\sum_{k=1}^{k_{1}}R_{u}[k]\geq\sum_{k=1}^{k_{1}}R^{\prime}_{u}[k]$.
Thus, the optimal trajectory should always move the MA towards $x^{*}$. Under this premise, if the trajectory $x^{\prime}[k],k\in{\cal K}$ moves the MA at a lower velocity than $v_{\max}$, it can also be seen that $R_{u}[k]>R^{\prime}_{u}[k]$. By combining the above results, we have $\frac{1}{K}\sum_{k=1}^{K}R_{u}[k]\geq\frac{1}{K}\sum_{k=1}^{K}R^{\prime}_{u}[k]$.
@@
Case 2: $x^{*}$ cannot be reached from $x_{0}$ within the time duration $T$, i.e., $\frac{\lvert x^{*}-x_{0}\rvert}{v_{\max}\tau}>K$. In this case, we can conduct a similar analysis as in Case 1 to show $\frac{1}{K}\sum_{k=1}^{K}R_{u}[k]\geq\frac{1}{K}\sum_{k=1}^{K}R^{\prime}_{u}[k]$. The details are omitted for brevity.

### 06 May 2026 v2

@@ III-A Two-Path Case
$\left|h[k]\right|^{2}=\left|a_{1}\right|^{2}+\left|a_{2}\right|^{2}+2\left|a_{1}\right|\left|a_{2}\right|\cos\varphi[k],$ (11)
@@ III-B Multi-Path Case
${W}_{i,j}=-\log_{2}\left(1+\frac{P_{t}\left|h\left(p_{i}\right)\right|^{2}}{\sigma^{2}}\right),\forall(i,j)\in\tilde{E}$, (15)
where $h\left(p_{i}\right)$ denotes the channel with the user when the MA is positioned at the $i$-th grid center. Based on the above, it can be shown that $\left(P1\right)$ is equivalent to the following fixed-hop shortest path problem $\left(P2\right)$.
@@ V Conclusion
Based on (16), the average rate $\frac{1}{K}\sum_{k=1}^{K}R[k]$ can be decomposed as $\frac{1}{K}\left(\sum_{k=1}^{k_{1}}R[k]+(K-k_{1})R[K]\right)$. Consider an arbitrary feasible trajectory, denoted as $x^{\prime}[k],k\in{\cal K}$, and let $R^{\prime}[k]$ denote the achievable rate at the receiver for $x=x^{\prime}[k]$. Evidently, we have $(K-k_{1})R[K]\geq\sum_{k=k_{1}}^{K}R^{\prime}[k]$. Next, we show $\sum_{k=1}^{k_{1}}R[k]\geq\sum_{k=1}^{k_{1}}R^{\prime}[k]$. Note that moving the MA from $x_{0}$ to $x^{\prime}$ within the first to the $k_{1}$-th time slot always yields an increasing channel power gain (see the green curve in Fig. 8), while moving the MA in the opposite direction always yields a decreasing channel power gain first within this period (see the red curve in Fig. 8). Hence, if the trajectory $x^{\prime}[k],k\in{\cal K}$ moves the MA away from $x^{*}$, we have $R[k]>R^{\prime}[k],1\leq k\leq k_{1}$, and thus, $\sum_{k=1}^{k_{1}}R[k]\geq\sum_{k=1}^{k_{1}}R^{\prime}[k]$.
Thus, the optimal trajectory should always move the MA towards $x^{*}$. Under this premise, if the trajectory $x^{\prime}[k],k\in{\cal K}$ moves the MA at a lower velocity than $v_{\max}$, it can also be seen that $R[k]>R^{\prime}[k]$. By combining the above results, we have $\frac{1}{K}\sum_{k=1}^{K}R[k]\geq\frac{1}{K}\sum_{k=1}^{K}R^{\prime}[k]$.
@@
Case 2: $x^{*}$ cannot be reached from $x_{0}$ within the time duration $T$, i.e., $\frac{\lvert x^{*}-x_{0}\rvert}{v_{\max}\tau}>K$. In this case, we can conduct a similar analysis as in Case 1 to show $\frac{1}{K}\sum_{k=1}^{K}R[k]\geq\frac{1}{K}\sum_{k=1}^{K}R^{\prime}[k]$. The details are omitted for brevity.